OpenStudy (anonymous):

How to solve and answer l 4x+3 l=5x+7

OpenStudy (anonymous):

the absolute value is the same thing, the only way it would change is if the equation inside was negative, so the first thing to do is subtract 5x from both sides.

OpenStudy (anonymous):

then you will have, -x + 3 = 7, subtract three then divide both sides by negative 1 and you get x = -4

OpenStudy (nurali):

l 4x+3 l=5x+7 4x+3=5x+7 or 4x+3=-5x-7 4x-5x=7-3 or 4x+5x=-7-3 -x=4 or 9x=-10 x=-4 or x=-10/9

OpenStudy (anonymous):

okay thanks I get :)

OpenStudy (anonymous):

I hope that helped!

OpenStudy (anonymous):

it did

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Please click best response.